please solve the logarithmic equation.
log(8)^x=(4/3)
so far i have.....x=8^(4/3)

Answer by jim_thompson5910(35256)

8^(4/3) = (2^3)^(4/3) = (cube root of 2^3)^4 = 2^4 = 16


So x = 16